OpenStudy (anonymous):

If 4x^2 + kx + 25 = 0 has exactly one solution, what could be the value of k?

OpenStudy (anonymous):

discriminant should be 0 k^2-4*4*25=0 k^2=400 k=+-20

OpenStudy (anonymous):

u can also cross check when k=20 then 4x^2 + kx + 25 = 0 reduces to (2x+5)^2 when k=-20 then 4x^2 + kx + 25 = 0 reduces to (2x-5)^2